# Non-stationary Gaussian random fields on hypersurfaces: Sampling and strong error analysis

### v1

@@ Abstract.
A flexible model for non-stationary Gaussian random fields on hypersurfaces is introduced. The class of random fields on curves and surfaces is characterized by a power spectral density of a second order elliptic differential operator. Sampling is done by a Galerkin–Chebyshev approximation based on the surface finite element method and Chebyshev polynomials. Strong error bounds are shown with convergence rates depending on the smoothness of the approximated random field.
@@
Acknowledgment: The authors thank Julie Rowlett and Iosif Polterovich for helpful discussions. This work was supported in part by the European Union (ERC, StochMan, 101088589), by the Mines Paris - PSL / INRAE “Geolearning” chair, by the Swedish Research Council (VR) through grant no. 2020-04170, by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Chalmers AI Research Centre (CHAIR)
@@ 1. Introduction
Random fields are powerful tools for modeling spatially dependent data. They have found uses in a wide range of applications, for instance in geostatistics, cosmological data analysis, climate modeling, and biomedical imaging (Marinucci and Peccati, 2011; Farag, 2014). One challenge in the modeling of spatial data is non-stationary behavior, i.e., different behaviors in different parts of the domain. Another challenge is that the domain may be a non-Euclidean space, for instance, a surface such as the sphere or on the cortical surface of the brain. In this paper, we present a surface finite element-based method to sample a flexible class of non-stationary random fields on curves and surfaces and show its strong convergence. The method, building on the foundational work for stationary fields introduced in Lang and Pereira (2023), is an extension of the stochastic partial differential equation (SPDE) approach pioneered by Whittle (1963) and popularized by Lindgren et al. (2011). The idea behind our method is to color white noise by applying a function of an elliptic differential operator $\mathcal{L}$. Formally, we study Gaussian random fields on curves and surfaces of the form
@@
where $\gamma$ is a function, called the power spectral density, $\mathcal{L}$ is an elliptic differential operator and $\mathcal{W}$ denotes white noise. By letting the coefficients of the differential operator vary over the domain, we can obtain local, non-stationary behaviors. If $1/\gamma$ is well-defined over $\amsmathbb{R}_{+}$, one may formally view $\mathcal{Z}$ as the solution to the stochastic partial differential equation
@@
The computational method we use to solve Equation 1, i.e., sample the random fields, is based on the surface finite element method (SFEM), a computational method pioneered by Dziuk (1988); Dziuk and Elliott (2013) and that has been used in the context of the generation of random fields in for instance Bonito et al. (2024); Jansson et al. (2022); Lang and Pereira (2023). Our main mathematical contribution is a strong convergence result. Using a functional calculus approach to the finite element discretization error, we obtain a strong rate of convergence of order $\mathcal{O}(K_{\alpha}(h)h^{\min\{\alpha-d/4;2\}})$, where $d=2$ for surfaces and $d=1$ for curves, and $K_{\alpha}(h)$ is a dimension-dependent logarithmic factor.
The SPDE approach to random fields and their approximation have been studied previously for both surfaces and Euclidean geometries, examples include Bonito et al. (2024); Borovitskiy et al. (2020); Bolin et al. (2020); Cox and Kirchner (2020); Jansson et al. (2022); Lang and Pereira (2023); Lindgren et al. (2011, 2022). However, to the best of our knowledge, we are the first to present a strong error analysis for the non-stationary case on hypersurfaces. Moreover, we achieve this without requiring an explicit approximation of the eigenfunctions of the elliptic operator, and contrary to for instance Borovitskiy et al. (2020), computing the eigenfunctions. In fact, the computation of eigenfunctions, with theoretical guarantees, is a notoriously difficult problem, see e.g., Boffi (2010). Our method circumvents this issue by introducing a Chebyshev approximation.
@@
The paper is structured as follows: In Section 2, we introduce the relevant deterministic framework. We provide the necessary background on geometry and functional analysis in Section 2.1. This is followed by a description of the main computational tool, surface finite elements in Section 2.2. Finally, Section 2.3 provides the relevant error estimates in the deterministic setting. In Section 3, we collect all material in the stochastic setting. We introduce first the class of considered random fields in Section 3.1 and their Galerkin–Chebyshev approximation in Section 3.2. The proof of its strong convergence is split into the SFEM error in Section 3.3 and the Chebyshev approximation error Section 3.4. For numerical confirmation of our theoretical results, we refer to our earlier work Lang and Pereira (2023) and the results presented in Bonito et al. (2024). The source code used to generate the figures is available at this address: https://github.com/mike-pereira/SFEMsim.
@@ 2.1. Geometric and functional analytic setting
where $k\in\amsmathbb{N}$ is the smallest integer such that $02k+s>$ 0. In this case, the corresponding norm is given by $\|u\|_{H^{s}(\mathcal{M})}=\|v\|_{H^{2k+s}(\mathcal{M})}$. We set $H^{0}(\mathcal{M})=L^{2}(\mathcal{M})$. The reader is referred to Herrmann et al. (2018), Strichartz (1983), Triebel (1985), and references therein for more details on Sobolev spaces defined using Bessel potentials.
@@
where for any $x_{0}\in\mathcal{M}$ the diffusion matrix $\mathcal{D}(x_{0})=[D_{ij}(x_{0})]_{i,j=1}^{d+1}$ is a real-valued, symmetric matrix such that for any $w\in T_{x_{0}}\mathcal{M}$, $\mathcal{D}(x_{0})w\in T_{x_{0}}\mathcal{M}$ and $(\mathcal{D}(x_{0})w)\cdot\overline{w}>0$ if $w\neq 0$. In particular, since $T_{x_{0}}\mathcal{M}=[\nu(x_{0})^{\perp}]$, $\mathcal{D}(x_{0})$ is simply a matrix admitting $\nu(x_{0})$ as an eigenvector with some eigenvalue $\mu_{1}(x_{0})\in\amsmathbb{R}$, and such that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 associated with its other eigenvectors are positive. Without loss of generality, we may assume that $\mu_{1}(x_{0})=0$, meaning that $\mathcal{D}\nu=0$, that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 are uniformly lower-bounded and upper-bounded on $\mathcal{M}$ by positive constants, and that $D_{ij}\in L^{\infty}(\mathcal{M})$ for any $11\leq i,j\leq d+$ 1. Finally, we assume that $V\in L^{\infty}(\mathcal{M})$ is a real-valued function that satisfies $V_{-}\leq V\leq V_{+}$ for some $0<V_{-}\leq V_{+}<+\infty$.
@@
Following Yagi (2010, Equation (1.33)), $\mathsf{A}_{\mathcal{M}}$ gives rise to an associated elliptic differential operator $\mathcal{L}:H^{1}(\mathcal{M})\rightarrow H^{-1}(\mathcal{M})$ defined weakly by
@@
Since the operator $\mathcal{L}$ differs from the Laplace–Beltrami operator only by a zeroth-order potential term and a diffusion function in the second order term, switching between the two operators corresponds to a change of metric on $\mathcal{M}$. Therefore, the eigenvalue problem for $\mathcal{L}$ is equivalent to that for the Laplace–Beltrami operator on $\mathcal{M}$ equipped with a possibly rough metric if the coefficients of $\mathcal{D}$ are not smooth. The results in Bandara et al. (2021) imply growth rates on the eigenvalues in accordance with Weyl’s law, and more specifically that there exist $c_{\mathcal{M}},C_{\mathcal{M}}>0$ such that for any $i\in\amsmathbb{N}$
@@
As a last step in this subsection, we introduce nonlinear functions of $\mathcal{L}$ which allow later in Section 3 for the definition of a variety of Gaussian random fields. For that, we call a function $\gamma:\amsmathbb{R}_{+}\rightarrow\amsmathbb{R}$ an $\alpha$-power spectral density if
@@
Applying a power spectral density to $\mathcal{L}$ results in a linear operator $\gamma(\mathcal{L})$ whose action on functions $f\in L^{2}(\mathcal{M})$ is defined by
@@
where $\{(\lambda_{i},e_{i})\}_{i\in\amsmathbb{N}}$ are the eigenpairs of $\mathcal{L}$ defined in Proposition 2.1. A typical example is the function $\gamma(\lambda)=(\kappa^{2}+\lambda)^{-\alpha}$ for $\alpha>d/4$ and $\kappa>0$, which can be used to obtain Whittle–Matérn random fields (Lang and Pereira, 2023).
The goal of the remainder of this section is to study the approximation of functions of the form $u=\gamma(\mathcal{L})f$, where $f\in L^{2}(\mathcal{M})$. Formally, if $1/\gamma$ is well-defined on the spectrum of $\mathcal{L}$, then this is the solution to the partial differential equation $\gamma(\mathcal{L})^{-1}u=f$.
@@ 2.2. SFEM–Galerkin approximation
The idea behind the surface finite element method, as introduced by Dziuk (1988), is to work on a polyhedral approximation of the surface that is in some sense close to the true surface $\mathcal{M}$. More precisely, fix $h>0$ and let $\mathcal{M}_{h}$ be a piecewise polygonal surface consisting of non-degenerate simplices (for $d=2$, triangles and for $d=1$, line segments) with vertices on $\mathcal{M}$, and such that $h$ is the size of the largest simplex defined as the in-ball radius. The set of simplices making up the discretized surface is denoted by $\mathcal{T}_{h}$, thus meaning that
@@
Following (Dziuk and Elliott, 2013, Section 1.4.1), we assume that the triangulation $\mathcal{T}_{h}$ is quasi-uniform, shape-regular, and that the number of simplices sharing the same vertex can be upper-bounded by a constant independent of $h$. In turn, these two assumption imply that $N_{h}\propto h^{-d}$ where $N_{h}\in\amsmathbb{N}$ denotes the number of vertices of $\mathcal{M}_{h}$.
@@
The discrete surface $\mathcal{M}_{h}$ is close to the true surface $\mathcal{M}$ in the sense that $\mathcal{M}_{h}$ is contained in a small neighborhood around $\mathcal{M}$ defined as follows. First, note that $\mathcal{M}$ can be seen as the boundary of some bounded open set $G\subset\amsmathbb{R}^{d+1}$ with exterior normal $\nu$. Then, following Dziuk and Elliott (2013, Section 2.3), we consider that there exists some (small) $\varpi>0$ such that $\mathcal{M}_{h}$ is contained in a so-called tubular neighborhood $U_{\varpi}$ of $\mathcal{M}$ defined by
@@
Following (Bonito et al., 2020, Section 1.2.1), we denote by $\sigma:\mathcal{M}_{h}\rightarrow\amsmathbb{R}^{+}$ the area element given by $\sigma=\mathrm{d}A/\mathrm{d}A_{h}$, such that for all $v\in L^{2}(\mathcal{M})$,
@@
The mapping $a$ is used to define the gradient of functions on $\mathcal{M}_{h}$ (Dziuk and Elliott, 2013). More specifically, for a differentiable $\eta:\mathcal{M}_{h}\to\amsmathbb{C}$, the gradient is given by
@@
(11) $\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}+\int_{\mathcal{M}_{h}}(V^{-\ell}u_{h})\overline{v}_{h}\,\mathrm{d}A_{h},\quad u_{h},v_{h}\in H^{1}(\mathcal{M}_{h}),$
where $\mathcal{D}^{-\ell}=[D_{ij}^{-\ell}]_{i,j=1}^{d+1}$. As in Dziuk and Elliott (2013), we assume that there exists $h_{0}\leavevmode\nobreak\ \in\leavevmode\nobreak\ (0,1)$ small enough, such that $\mathsf{A}_{\mathcal{M}_{h}}$ is coercive (and continuous) whenever $h\leq h_{0}$. Unless stated otherwise, we now assume that this last condition on $h$ is fulfilled.
To conclude this subsection, we introduce the (linear) finite element space $S_{h}$ on $\mathcal{M}_{h}$. The finite element space $S_{h}$ is defined as the complex span of the standard real-valued nodal basis $\psi_{1},\dots,\psi_{N_{h}}\colon\mathcal{M}\rightarrow\amsmathbb{R}$, where for any $i\in\{1,\dots,N_{h}\}$, $\psi_{i}|_{T}$ is a polynomial of at most degree one taking the value $1$ at the $i$-th vertex of $\mathcal{M}_{h}$ and 00 at all the other vertices, i.e.,
@@
Since the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$ are coercive, positive definite, Hermitian and have real coefficients, these two operators are diagonalizable in the sense that they each give rise to a set of $N_{h}$ eigenpairs (Hall, 2013). On the one hand, there exists a sequence $0\leq\Lambda_{1}^{h}\leq\dots\leq\Lambda_{N_{h}}^{h}$ and an $L^{2}(\mathcal{M}_{h})$-orthonormal basis $E_{1}^{h},\ldots,E_{N_{h}}^{h}$ of $S_{h}$ such that
@@
In particular, using the same approach as in Proposition 2.1, we can assume that the eigenfunctions $\{E_{i}^{h}\}_{1\leq i\leq N_{h}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are all real-valued. The eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$ are linked to one another through the following lemma, due to Bonito et al. (2018, Lemma 3.1), Bonito et al. (2024, Lemma 4.1) and Strang and Fix (2008, Theorem 6.1).
@@
As defined, $\bm{C}$ is a symmetric positive definite matrix and $\bm{R}$ is a symmetric positive semi-definite matrix. Let then $\sqrt{\bm{C}}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ be an invertible matrix satisfying $\sqrt{\bm{C}}(\sqrt{\bm{C}})^{T}=\bm{C}$. Then, by Lang and Pereira (2023, Corollary 3.2), the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ are also the eigenvalues of the matrix $\bm{S}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ defined by

### v2

@@ Abstract.
A flexible model for non-stationary Gaussian random fields on hypersurfaces is introduced. The class of random fields on curves and surfaces is characterized by an amplitude spectral density of a second order elliptic differential operator. Sampling is done by a Galerkin–Chebyshev approximation based on the surface finite element method and Chebyshev polynomials. Strong error bounds are shown with convergence rates depending on the smoothness of the approximated random field. Numerical experiments that confirm the convergence rates are presented.
@@
Acknowledgment: The authors thank Mihály Kovács, Julie Rowlett and Iosif Polterovich for helpful discussions. This work was supported in part by the European Union (ERC, StochMan, 101088589), by the Mines Paris - PSL / INRAE “Geolearning” chair, by the Swedish Research Council (VR) through grant no. 2020-04170, by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Chalmers AI Research Centre (CHAIR). Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1. Introduction
Random fields are powerful tools for modeling spatially dependent data. They have found uses in a wide range of applications, for instance in geostatistics, cosmological data analysis, climate modeling, and biomedical imaging [22, 12]. One challenge in the modeling of spatial data is non-stationary behavior, i.e., different behaviors in different parts of the domain. Another challenge is that the domain may be a non-Euclidean space, for instance, a surface such as the sphere or on the cortical surface of the brain. In this paper, we present a surface finite element-based method to sample a flexible class of non-stationary random fields on curves and surfaces and show its strong convergence. The method, building on the foundational work for stationary fields introduced in [19], is an extension of the stochastic partial differential equation (SPDE) approach pioneered by [27] and popularized by [21]. The idea behind our method is to color white noise by applying a function of an elliptic differential operator $\mathcal{L}$. Formally, we study Gaussian random fields on curves and surfaces of the form
@@
where $\mathcal{L}$ is an elliptic differential operator, $\mathcal{W}$ denotes white noise, $\gamma$ is a function, called amplitude spectral density in analogy to the spectral analysis of time signals [15]. By letting the coefficients of the differential operator vary over the domain, we can obtain local, non-stationary behaviors. If $1/\gamma$ is well-defined over $\amsmathbb{R}_{+}$, one may formally view $\mathcal{Z}$ as the solution to the stochastic partial differential equation
@@
The computational method we use to solve Equation 1, i.e., sample the random fields, is based on the surface finite element method (SFEM), a computational method pioneered by [10, 11] and that has been used in the context of the generation of Whittle–Matérn random fields with Laplace–Beltrami operators in for instance [6, 18, 19]. Our main mathematical contribution is a strong convergence result for equations with more general elliptic operators and amplitude spectral densities than fractional powers. Using a functional calculus approach to the finite element discretization error, we obtain a strong rate of convergence of order $\mathcal{O}(C_{\alpha}(h)h^{2\min\{\alpha-d/4;1\}})$, where $d=2$ for surfaces and $d=1$ for curves, and $C_{\alpha}(h)$ is a dimension-dependent logarithmic factor.
The SPDE approach to random fields and their approximation have been studied previously for both surfaces and Euclidean geometries, examples include [6, 8, 3, 9, 18, 19, 21, 20]. However, to the best of our knowledge, we are the first to present a strong error analysis for the non-stationary case on hypersurfaces that are only given by mesh points. In contrast to the method presented in [16], we do not need to explicitly know the true hypersurface and perform computations on it. Instead, we only require a polyhedral approximation of the hypersurface. Moreover, like in [9], we achieve this without requiring explicit approximation bounds on the eigenfunctions of the elliptic operator, and contrary to for instance [8], computing the eigenfunctions. In fact, the computation of eigenfunctions, with theoretical guarantees, is a notoriously difficult problem, see e.g., [2]. Our method circumvents this issue by introducing a Chebyshev approximation.
@@
The paper is structured as follows: In Section 2, we introduce the relevant deterministic framework. We provide the necessary background on geometry and functional analysis in Section 2.1. This is followed by a description of the main computational tool, surface finite elements in Section 2.2. Finally, Section 2.3 provides the relevant error estimates in the deterministic setting. In Section 3, we collect all material in the stochastic setting. We introduce first the class of considered random fields in Section 3.1 and their Galerkin–Chebyshev approximation in Section 3.2. The proof of its strong convergence is split into the SFEM error in Section 3.3 and the Chebyshev approximation error in Section 3.4. In Section 4, we present numerical experiments that confirm the strong error bounds. The source code used to generate the figures is available at this address: https://github.com/mike-pereira/SFEMsim.
@@ 2.1. Geometric and functional analytic setting
where $k\in\amsmathbb{N}$ is the smallest integer such that $02k+s>$ 0. In this case, the corresponding norm is given by $\|u\|_{H^{s}(\mathcal{M})}=\|v\|_{H^{2k+s}(\mathcal{M})}$. We set $H^{0}(\mathcal{M})=L^{2}(\mathcal{M})$. The reader is referred to [17], [25], [26], and references therein for more details on Sobolev spaces defined using Bessel potentials.
@@
where for any $x_{0}\in\mathcal{M}$ the diffusion matrix $\mathcal{D}(x_{0})=[D_{ij}(x_{0})]_{i,j=1}^{d+1}$ is a real-valued, symmetric matrix such that for any $w\in T_{x_{0}}\mathcal{M}$, $\mathcal{D}(x_{0})w\in T_{x_{0}}\mathcal{M}$ and $(\mathcal{D}(x_{0})w)\cdot\overline{w}>0$ if $w\neq 0$. In particular, since $T_{x_{0}}\mathcal{M}=[\nu(x_{0})^{\perp}]$, $\mathcal{D}(x_{0})$ is simply a matrix admitting $\nu(x_{0})$ as an eigenvector with some eigenvalue $\mu_{1}(x_{0})\in\amsmathbb{R}$, and such that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 associated with its other eigenvectors are positive. Without loss of generality, we may assume that $\mu_{1}(x_{0})=0$, meaning that $\mathcal{D}\nu=0$, that the eigenvalues $\mu_{i}(x_{0})$, $12\leq i\leq d+$ 1 are uniformly lower-bounded and upper-bounded on $\mathcal{M}$ by positive constants, and for simplicity that $D_{ij}\in C^{\infty}(\mathcal{M})$ for any $11\leq i,j\leq d+$ 1. Finally, we assume that $V\in L^{\infty}(\mathcal{M})$ is a real-valued function that satisfies $V_{-}\leq V\leq V_{+}$ for some $0<V_{-}\leq V_{+}<+\infty$.
@@
Following [28, Equation (1.33)], $\mathsf{A}_{\mathcal{M}}$ gives rise to an associated elliptic differential operator $\mathcal{L}:H^{1}(\mathcal{M})\rightarrow H^{-1}(\mathcal{M})$ defined weakly by
@@
Since the operator $\mathcal{L}$ differs from the Laplace–Beltrami operator only by a zeroth-order potential term and a diffusion function in the second order term, switching between the two operators corresponds to a change of metric on $\mathcal{M}$. Therefore, the eigenvalue problem for $\mathcal{L}$ is equivalent to that for the Laplace–Beltrami operator on $\mathcal{M}$ equipped with a possibly rough metric if the coefficients of $\mathcal{D}$ are not smooth. The results in [1] imply growth rates on the eigenvalues in accordance with Weyl’s law, and more specifically that there exist $c_{\mathcal{M}},C_{\mathcal{M}}>0$ such that for any $i\in\amsmathbb{N}$
@@
As a last step in this subsection, we introduce nonlinear functions of $\mathcal{L}$ which allow later in Section 3 for the definition of a variety of Gaussian random fields. For that, we call a function $\gamma:\amsmathbb{R}_{+}\rightarrow\amsmathbb{R}$ an $\alpha$-amplitude spectral density if
@@
Applying a amplitude spectral density to $\mathcal{L}$ results in a linear operator $\gamma(\mathcal{L})$ whose action on functions $f\in L^{2}(\mathcal{M})$ is defined by
@@
where $\{(\lambda_{i},e_{i})\}_{i\in\amsmathbb{N}}$ are the eigenpairs of $\mathcal{L}$ defined in Proposition 2.1. A typical example is the function $\gamma(\lambda)=(\kappa^{2}+\lambda)^{-\alpha}$ for $\alpha>d/4$ and $\kappa>0$, which can be used to obtain Whittle–Matérn random fields [19].
Remark 2.2.
For any $\alpha$-amplitude spectral density $\gamma$ and any $f\in L^{2}(\mathcal{M})$, $\gamma(\mathcal{L})f\in L^{2}(\mathcal{M})$. In fact, we have for any $s\in[0,\alpha]$, $\|\mathcal{L}^{s}(\gamma(\mathcal{L})f)\|_{L^{2}(\mathcal{M})}<\infty$. Indeed,
$\|\mathcal{L}^{s}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}=\sum_{i=1}^{\infty}|\lambda_{i}^{s}\gamma(\lambda_{i})|^{2}|(f,e_{i})_{L^{2}(\mathcal{M})}|^{2},$
where, since $\gamma$ is an $\alpha$-amplitude spectral density and $\lambda_{i}\in H_{\pi/2}$ for any $i\in\amsmathbb{N}$, $|\lambda_{i}^{s}\gamma(\lambda_{i})|\lesssim|\lambda_{i}|^{-(\alpha-s)}$. Using then the fact that for any $i\in\amsmathbb{N}$, $\lambda_{i}\geq\lambda_{1}>0$, and that $\alpha-s\geq 0$, we conclude that $|\lambda_{i}^{s}\gamma(\lambda_{i})|\lesssim|\lambda_{1}|^{-(\alpha-s)}\lesssim 1$ and therefore $\|\mathcal{L}^{s}\gamma(\mathcal{L})f\|_{L^{2}(\mathcal{M})}\lesssim\sum_{i=1}^{\infty}|(f,e_{i})_{L^{2}(\mathcal{M})}|^{2}=\|f\|_{L^{2}(\mathcal{M})}<\infty$.
The goal of the remainder of this section is to study the approximation of functions of the form $u=\gamma(\mathcal{L})f$, where $f\in L^{2}(\mathcal{M})$. Formally, if $1/\gamma$ is well-defined on the spectrum of $\mathcal{L}$, then this is the solution to the partial differential equation $(1/\gamma)(\mathcal{L})u=f$.
@@ 2.2. SFEM–Galerkin approximation
The idea behind the surface finite element method, as introduced by [10], is to work on a polyhedral approximation of the surface that is in some sense close to the true surface $\mathcal{M}$. More precisely, fix $h>0$ and let $\mathcal{M}_{h}$ be a piecewise polygonal surface consisting of non-degenerate simplices (for $d=2$, triangles and for $d=1$, line segments) with vertices on $\mathcal{M}$, and such that $h$ is the size of the largest simplex defined as the in-ball radius. The set of simplices making up the discretized surface is denoted by $\mathcal{T}_{h}$, thus meaning that
@@
Following [11, Section 1.4.1], we assume that the triangulation $\mathcal{T}_{h}$ is quasi-uniform, shape-regular, and that the number of simplices sharing the same vertex can be upper-bounded by a constant independent of $h$. In turn, these two assumption imply that $N_{h}\propto h^{-d}$ where $N_{h}\in\amsmathbb{N}$ denotes the number of vertices of $\mathcal{M}_{h}$.
@@
The discrete surface $\mathcal{M}_{h}$ is close to the true surface $\mathcal{M}$ in the sense that $\mathcal{M}_{h}$ is contained in a small neighborhood around $\mathcal{M}$ defined as follows. First, note that $\mathcal{M}$ can be seen as the boundary of some bounded open set $G\subset\amsmathbb{R}^{d+1}$ with exterior normal $\nu$. Then, following [11, Section 2.3], we consider that there exists some (small) $\varpi>0$ such that $\mathcal{M}_{h}$ is contained in a so-called tubular neighborhood $U_{\varpi}$ of $\mathcal{M}$ defined by
@@
Following [4, Section 1.2.1], we denote by $\sigma:\mathcal{M}_{h}\rightarrow\amsmathbb{R}^{+}$ the area element given by $\sigma=\mathrm{d}A/\mathrm{d}A_{h}$, such that for all $v\in L^{2}(\mathcal{M})$,
@@
The mapping $a$ is used to define the gradient of functions on $\mathcal{M}_{h}$ [11]. More specifically, for a differentiable $\eta:\mathcal{M}_{h}\to\amsmathbb{C}$, the gradient is given by
@@
(11) $\mathsf{A}_{\mathcal{M}_{h}}(u_{h},v_{h})=\int_{\mathcal{M}_{h}}(\mathcal{D}^{-\ell}\nabla_{\mathcal{M}_{h}}u_{h})\cdot(\nabla_{\mathcal{M}_{h}}\overline{v}_{h})\,\mathrm{d}A_{h}+\int_{\mathcal{M}_{h}}(V^{-\ell}u_{h})\overline{v}_{h}\,\mathrm{d}A_{h},$
where $u_{h},v_{h}\in H^{1}(\mathcal{M}_{h})$ and $\mathcal{D}^{-\ell}=[D_{ij}^{-\ell}]_{i,j=1}^{d+1}$. As in [11], we assume that there exists $h_{0}\leavevmode\nobreak\ \in\leavevmode\nobreak\ (0,1)$ small enough, such that $\mathsf{A}_{\mathcal{M}_{h}}$ is coercive (and continuous) whenever $h\leq h_{0}$. Unless stated otherwise, we now assume that this last condition on $h$ is fulfilled.
To conclude this subsection, we introduce the (linear) finite element space $S_{h}$ on $\mathcal{M}_{h}$. The finite element space $S_{h}$ is defined as the complex span of the standard real-valued nodal basis
$\psi_{1},\dots,\psi_{N_{h}}\colon\mathcal{M}\rightarrow\amsmathbb{R},$
where for any $i\in\{1,\dots,N_{h}\}$, $\psi_{i}|_{T}$ is a polynomial of at most degree one taking the value $1$ at the $i$-th vertex of $\mathcal{M}_{h}$ and 00 at all the other vertices, i.e.,
@@
Since the bilinear forms $\mathsf{A}_{\mathcal{M}}$ and $\mathsf{A}_{\mathcal{M}_{h}}$ are coercive, positive definite, Hermitian and have real coefficients, these two operators are diagonalizable in the sense that they each give rise to a set of $N_{h}$ eigenpairs [14]. On the one hand, there exists a sequence $0\leq\Lambda_{1}^{h}\leq\dots\leq\Lambda_{N_{h}}^{h}$ and an $L^{2}(\mathcal{M}_{h})$-orthonormal basis $E_{1}^{h},\ldots,E_{N_{h}}^{h}$ of $S_{h}$ such that
@@
In particular, using the same approach as in Proposition 2.1, we can assume that the eigenfunctions $\{E_{i}^{h}\}_{1\leq i\leq N_{h}}$ and $\{e_{i}^{h}\}_{1\leq i\leq N_{h}}$ are all real-valued. The eigenvalues of the operators $\mathcal{L}$, $\mathcal{L}_{h}$ and $\mathsf{L}_{h}$ are linked to one another through the following lemma, due to [5, Lemma 3.1], [6, Lemma 4.1] and [24, Theorem 6.1].
@@
As defined, $\bm{C}$ is a symmetric positive definite matrix and $\bm{R}$ is a symmetric positive semi-definite matrix. Let then $\sqrt{\bm{C}}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ be an invertible matrix satisfying $\sqrt{\bm{C}}(\sqrt{\bm{C}})^{T}=\bm{C}$. Then, by [19, Corollary 3.2], the eigenvalues $\{\Lambda_{i}^{h}\}_{1\leq i\leq N_{h}}$ are also the eigenvalues of the matrix $\bm{S}\in\amsmathbb{R}^{N_{h}\times N_{h}}$ defined by